# ZEUS: Zero-shot Embeddings for Unsupervised Separation of Tabular Data

### v1

@@ Appendix C Pre-training plots
Figure 4 presents two plots of average ARI over 200 synthetic validation datasets throughout 300 pre-training epochs. The first plot corresponds to the Gaussian-categorical datasets referred to as Synthetic Gaussian, while the second illustrates their NN transformed variants, named Synthetic Transformed. The data generation procedures are described in more detail in Appendix B. Both plots illustrate that the quality of the ZEUS representation improves over time, indicating that the model is learning new patterns, as evidenced by the increasing ARI during the pre-training process.
@@ Appendix D Statistics of datasets used in experimental study
Tables 5, 6, and 7 provide detailed information about the datasets used in the experimental analysis. Each table includes the number of set instances, the number of categorical and numerical features, the total dimensionality, and the overall number of categories, represented by the length of the one-hot encoded vectors. Additionally, the tables report the number of classes that each dataset contains. In Table 5, the ID column corresponds to the OpenML ID, while in the remaining tables it functions solely as an index.
@@ Appendix E Baselines
    scikit-learn - used for k-means and GMM
@@
    https://github.com/vlukiyanov/pt-dec - implementation of the DEC
@@
    https://github.com/dawnranger/IDEC-pytorch - source code for the IDEC
@@
Most parameters of the k-means and GMM methods are left at their default values. Only the $n\_init$ option was increased to 100 for k-means and 50 for GMM in order to improve stability.
@@
As mentioned in Section 3.1, a standard autoencoder with hidden layers $[500,500,2000]$ is used for each AE-based method. The resulting network is first pre-trained for 1000 epochs and then fine-tuned for up to 1000 additional epochs, separately for each dataset and model. Following the GCEALS evaluation procedure, multiple latent dimension sizes $[5,10,15,20]$ were tested across all considered methods. The results indicate that the default latent dimension of 10 does not yield the best performance; instead, a dimension of 20 generally performs better. Changing other hyperparameters, including the learning rate, optimizer, and clustering loss weight $\gamma$, generally did not lead to improved results. Therefore, the remaining parameters were left at their default values as proposed by the authors of the respective repositories.
Table 8: Real/OpenML datasets - ARI
ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
14 38.74 45.90 40.94 45.36 49.60 36.78 45.04 50.56
15 82.85 71.38 61.44 71.74 86.29 87.48 24.38 81.28
16 55.62 64.75 61.66 69.97 68.56 55.61 55.27 74.03
18 54.86 46.88 51.45 51.47 50.00 53.54 49.93 51.63
22 35.48 47.39 31.09 50.80 50.02 35.69 28.38 56.05
35 70.68 70.68 70.22 70.48 68.63 64.58 56.57 85.12
51 28.35 -2.54 37.82 34.13 38.13 34.58 35.13 43.66
53 45.21 5.17 39.14 40.62 43.50 41.91 21.32 35.76
56 57.79 58.49 57.10 58.31 56.97 55.20 55.66 66.41
61 62.01 90.39 59.79 60.21 57.84 60.14 51.37 85.15
187 89.75 93.09 82.87 84.86 85.02 85.10 55.67 88.19
377 56.70 58.47 62.06 59.52 63.03 59.01 59.79 55.30
458 95.08 96.97 64.33 75.70 94.58 63.29 68.31 99.19
481 58.51 2.77 36.84 48.03 57.28 52.93 19.67 8.68
694 35.91 43.30 35.26 36.56 26.30 36.02 30.76 33.38
721 43.28 43.28 13.12 25.12 17.75 28.77 -0.03 43.28
733 50.97 46.29 38.39 55.11 67.47 73.51 26.26 74.97
745 57.63 55.44 78.35 75.52 46.37 50.94 57.81 73.85
756 55.70 4.46 72.81 53.40 58.60 65.58 44.14 50.31
796 49.62 2.26 76.54 52.19 50.07 55.92 43.40 13.99
820 50.16 36.77 44.29 49.24 51.21 49.30 24.51 28.87
840 39.05 48.37 53.60 46.72 49.14 39.16 27.95 27.66
854 76.14 -0.21 62.33 60.83 35.05 70.52 72.25 76.14
1462 1.32 0.31 10.03 10.17 1.00 3.80 8.48 92.28
1495 96.81 98.40 72.74 82.80 96.81 91.28 77.51 7.35
1499 77.33 62.99 73.68 75.28 69.99 78.48 29.60 82.40
1510 67.07 78.02 63.73 59.77 71.32 69.84 69.47 74.26
1523 21.16 41.21 25.48 24.57 28.16 23.05 18.50 19.60
4153 55.78 58.03 59.15 46.23 56.41 58.44 39.43 62.29
40496 54.02 34.79 42.35 37.51 49.74 42.35 37.33 32.25
40682 58.32 86.29 57.89 59.22 57.39 86.05 40.58 53.15
40705 43.49 37.92 0.73 30.63 47.90 46.28 10.71 45.45
42261 62.01 90.39 60.13 58.15 59.78 56.78 49.09 85.15
42585 60.84 30.52 51.26 60.71 91.79 43.40 38.30 95.05
Mean 55.54 48.49 51.43 53.56 55.93 54.57 40.37 57.43
Mean-rank 3.90 4.57 4.72 4.29 3.96 4.31 6.65 3.60
Top-3 16 14 9 8 18 13 2 21
Top-1 4 8 4 0 3 1 0 12
Table 9: Synthetic gaussian datasets - ARI
ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
0 86.16 82.13 83.44 79.66 87.06 77.26 58.57 87.65
1 88.66 87.36 81.13 79.61 85.84 75.62 54.10 72.04
2 81.30 26.46 71.57 78.04 86.65 78.91 48.59 82.96
3 95.02 36.77 92.03 76.44 92.74 89.81 62.67 95.76
4 90.64 92.97 82.65 77.80 91.61 84.51 69.70 92.85
5 89.83 25.85 86.52 69.06 90.01 88.31 67.70 90.72
6 96.21 98.07 85.99 94.26 94.24 92.39 73.15 97.89
7 92.59 30.52 81.73 79.06 91.35 89.80 58.10 93.60
8 88.29 89.62 81.15 76.53 87.56 78.37 50.30 84.00
9 93.36 98.54 68.58 85.42 91.14 85.29 56.89 94.55
10 79.26 93.32 72.75 90.90 72.54 73.51 48.47 87.32
11 93.60 98.48 90.77 89.11 95.03 92.19 64.24 97.45
12 86.82 77.45 73.64 71.40 86.04 77.93 58.45 86.04
13 96.77 98.07 94.05 90.15 96.11 85.66 78.20 96.17
14 90.73 43.20 79.65 69.09 92.68 77.93 67.84 94.60
15 93.58 96.23 85.74 89.30 94.43 89.41 72.88 96.68
16 94.70 92.45 90.98 89.41 94.53 88.14 85.86 87.19
17 85.26 87.10 62.23 79.52 84.08 75.00 42.55 86.29
18 88.77 94.83 84.47 84.74 87.39 83.46 67.96 83.01
19 86.54 89.11 76.18 78.46 86.06 67.99 70.50 73.85
Mean 89.90 76.93 81.26 81.40 89.36 82.57 62.84 89.03
Mean-Rank 2.60 3.35 5.40 5.45 3.10 5.50 7.70 2.90
Top-3 16 13 0 1 15 0 0 15
Top-1 3 10 0 0 1 0 0 6
Table 10: Synthetic transformed datasets - ARI
ID KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
0 60.56 79.33 46.31 68.98 76.93 46.35 29.42 87.42
1 93.30 96.55 76.57 83.99 91.29 70.91 73.09 89.40
2 89.76 96.31 57.79 65.99 94.42 66.34 48.34 95.88
3 70.16 50.38 56.23 76.06 83.06 60.09 54.54 80.98
4 44.32 80.63 47.60 44.77 68.39 50.12 49.37 73.66
5 91.49 96.51 74.48 89.11 86.46 77.84 66.47 81.63
6 60.98 81.63 61.97 76.29 75.58 61.89 69.84 87.02
7 76.80 83.98 71.37 77.69 86.56 77.49 56.51 89.57
8 89.77 77.10 85.53 84.42 93.33 92.96 78.87 98.11
9 84.40 98.48 82.64 93.03 97.14 81.02 65.62 97.20
10 82.47 91.91 61.00 86.16 87.58 46.08 39.27 86.42
11 90.77 99.32 81.41 91.18 94.78 88.12 70.29 93.19
12 67.64 91.80 52.87 63.56 85.45 60.61 55.59 87.75
13 85.99 50.80 45.48 35.34 29.04 12.91 32.89 91.60
14 93.52 54.56 15.81 22.79 94.12 38.21 15.10 98.49
15 36.62 30.93 37.71 43.70 42.78 40.65 29.70 28.96
16 95.62 65.11 90.13 94.97 97.47 95.32 98.27 99.01
17 60.54 95.74 56.24 91.71 78.72 59.11 13.29 96.44
18 63.92 17.39 58.47 70.21 54.61 47.93 10.01 79.70
19 62.12 79.06 49.36 65.77 81.07 51.19 26.84 84.13
Mean 75.04 75.88 60.45 71.29 79.94 61.26 49.17 86.33
Mean-Rank 4.45 3.20 6.20 4.25 3.00 5.70 7.05 2.15
Top-3 5 14 0 6 15 2 1 17
Top-1 0 8 0 1 1 0 0 10
@@ F.1 How effective is ZEUS for clustering?
Tables 8, 9, and 10 contain the complete results for individual datasets from the experiments discussed in Section 3.2 in Tables 1 and 2, as presented in the main text. The reported values represent averages of ARI over 5 random seeds. In addition to the rows presenting outcomes for specific datasets, the tables include summary rows labeled Mean, Mean-Rank, Top-3, and Top-1. These represent, respectively: the average ARI, the average rank computed across all datasets, the number of times a given model appears in the top 3, and the total number of wins achieved by each method. It is worth noting that the Top-1 and Top-3 rows indicate clear wins and clear appearances in top-3 positions.
@@
The overall conclusions align with those presented in the primary analysis. One noteworthy point is that ZEUS achieves strong performance in terms of the Top-3 and Top-1 statistics, being outperformed in this regard only in Table 9, where k-means and GMM score better. Another interesting observation is that GMM struggles with categorical features, which significantly worsens its average scores in Tables 9 and 10.
@@ F.2 Are ZEUS’s assignments well calibrated?
Table 11, 12 and 13 contain detailed extension of Table 3
Table 11: Real/OpenML datasets - Brier
@@
Table 12: Synthetic gaussian datasets - Brier
@@
Table 13: Synthetic transformed datasets - Brier

### v2

@@ Appendix C Pre-training plots
Figure 4 presents two plots of average ARI over 200 synthetic validation datasets throughout 300 pre-training epochs. The first plot corresponds to the Gaussian-categorical datasets referred to as Synthetic Gaussian, while the second illustrates their NN transformed variants, named Synthetic Transformed. The data generation procedures are described in more detail in Appendix B.2. Both plots illustrate that the quality of the ZEUS representation improves over time, indicating that the model is learning new patterns, as evidenced by the increasing ARI during the pre-training process.
@@ Appendix D Statistics of datasets used in experimental study
Tables 6, 7, and 8 provide detailed information about the datasets used in the experimental analysis. Each table includes the number of set instances, the number of categorical and numerical features, the total dimensionality, and the overall number of categories, represented by the length of the one-hot encoded vectors. Additionally, the tables report the number of classes that each dataset contains. In Table 6, the ID column corresponds to the OpenML ID, while in the remaining tables it functions solely as an index.
@@ Appendix E Baselines
    scikit-learn - used for k-means and GMM,
@@
    https://github.com/vlukiyanov/pt-dec - implementation of the DEC,
@@
    https://github.com/dawnranger/IDEC-pytorch - source code for the IDEC,
@@
    https://github.com/jsvir/idc - official implementation of the IDC method,
  5. 5.
  6. 6.
    https://github.com/PriorLabs/tabpfn-extensions - a library that extends TabPFN functionality to a wide spectrum of machine learning tasks, including unsupervised ones,
  7. 7.
    https://github.com/clabrugere/pytorch-scarf - code repository for SCARF.
@@
Most parameters of the k-means and GMM methods are left at their default values. Only the n_initn\_init option was increased to 100 for k-means and 50 for GMM in order to improve stability.
@@
As mentioned in Section 3.1, a standard autoencoder with hidden layers $[500,500,2000]$ is used for each AE-based method. The resulting network is first pre-trained for 1000 epochs and then fine-tuned for up to 1000 additional epochs, separately for each dataset and model. Following the GCEALS evaluation procedure, multiple latent dimension sizes $[5,10,15,20]$ were tested across all considered methods. The results indicate that the default latent dimension of 10 does not yield the best performance; instead, a dimension of 20 generally performs better. Changing other hyperparameters, including the learning rate, optimizer, and clustering loss weight $\gamma$, generally did not lead to improved results. Therefore, the remaining parameters were left at their default values as proposed by the authors of the respective codebases.
For TabPFN, we use the get_embeddings_per_column method from the TabPFNUnsupervisedModel class, provided in the TabPFN Extensions repository [^1] https://github.com/PriorLabs/tabpfn-extensions, to extract per-column representations. These embeddings are then averaged across columns to obtain the final representation for each data point. For SCARF, we follow the procedure outlined in the example.ipynb notebook from the original repository [^2] https://github.com/clabrugere/pytorch-scarf. In both cases, we cluster the learned representations using $k$-means, consistent with our approach in ZEUS.
@@ F.1 How effective is ZEUS for clustering?
Tables 9, 10, and 11 contain the complete results for individual datasets from the experiments discussed in Section 3.2 in Tables 1 and 2, as presented in the main text. The reported values represent averages of ARI over 5 random seeds. In addition to the rows presenting outcomes for specific datasets, the tables include summary rows labeled Mean, Mean-Rank, Top-3, and Top-1. These represent, respectively: the average ARI, the average rank computed across all datasets, the number of times a given model appears in the top 3, and the total number of wins achieved by each method. It is worth noting that the Top-1 and Top-3 rows indicate clear wins and clear appearances in top-3 positions.
@@
The overall conclusions align with those presented in the primary analysis. One noteworthy point is that ZEUS achieves strong performance in terms of the Top-3 and Top-1 statistics, being outperformed in this regard only in Table 10, where k-means and GMM score better. Another interesting observation is that GMM struggles with categorical features, which significantly worsens its average scores in Tables 10 and 11.
Table 9: Comparison of clustering quality using ARI metric on OpenML datasets (higher score indicate better performance).
ID KM GMM AE-KM AE-GMM DEC IDEC IDC G-CEALS TabPFN SCARF ZEUS
14 38.74 45.90 40.94 45.36 49.60 36.78 45.03 45.04 22.02 1.90 50.56
15 82.85 71.38 61.44 71.74 86.29 87.48 83.20 24.38 74.31 75.01 81.28
16 55.62 64.75 61.66 69.97 68.56 55.61 65.84 55.27 14.69 3.74 74.03
18 54.86 46.88 51.45 51.47 50.00 53.54 44.40 49.93 51.50 5.02 51.63
22 35.48 47.39 31.09 50.80 50.02 35.69 46.03 28.38 11.09 2.63 56.05
35 70.68 70.68 70.22 70.48 68.63 64.58 77.99 56.57 52.24 74.61 85.12
51 28.35 -2.54 37.82 34.13 38.13 34.58 38.93 35.13 38.26 -1.27 43.66
53 45.21 5.17 39.14 40.62 43.50 41.91 33.03 21.32 26.61 24.24 35.76
56 57.79 58.49 57.10 58.31 56.97 55.20 59.25 55.66 58.42 61.26 66.41
61 62.01 90.39 59.79 60.21 57.84 60.14 58.30 51.37 77.20 51.19 85.15
187 89.75 93.09 82.87 84.86 85.02 85.10 80.04 55.67 37.55 40.05 88.19
377 56.70 58.47 62.06 59.52 63.03 59.01 62.39 59.79 66.89 14.45 55.30
458 95.08 96.97 64.33 75.70 94.58 63.29 86.77 68.31 61.41 2.87 99.19
481 58.51 2.77 36.84 48.03 57.28 52.93 43.17 19.67 2.77 1.79 8.68
694 35.91 43.30 35.26 36.56 26.30 36.02 33.47 30.76 27.68 14.09 33.38
721 43.28 43.28 13.12 25.12 17.75 28.77 12.74 -0.03 23.63 6.35 43.28
733 50.97 46.29 38.39 55.11 67.47 73.51 68.88 26.26 -8.35 49.41 74.97
745 57.63 55.44 78.35 75.52 46.37 50.94 59.29 57.81 1.26 16.75 73.85
756 55.70 4.46 72.81 53.40 58.60 65.58 53.32 44.14 10.60 64.82 50.31
796 49.62 2.26 76.54 52.19 50.07 55.92 79.59 43.40 0.88 58.25 13.99
820 50.16 36.77 44.29 49.24 51.21 49.30 44.59 24.51 50.26 13.04 28.87
840 39.05 48.37 53.60 46.72 49.14 39.16 19.59 27.95 8.89 32.13 27.66
854 76.14 -0.21 62.33 60.83 35.05 70.52 64.81 72.25 -0.54 6.48 76.14
1462 1.32 0.31 10.03 10.17 1.00 3.80 8.83 8.48 3.31 -0.03 92.28
1495 96.81 98.40 72.74 82.80 96.81 91.28 83.96 77.51 47.11 5.10 7.35
1499 77.33 62.99 73.68 75.28 69.99 78.48 67.82 29.60 5.82 24.93 82.40
1510 67.07 78.02 63.73 59.77 71.32 69.84 70.55 69.47 54.59 63.57 74.26
1523 21.16 41.21 25.48 24.57 28.16 23.05 31.35 18.50 2.45 21.04 19.60
4153 55.78 58.03 59.15 46.23 56.41 58.44 50.44 39.43 27.80 36.21 62.29
40496 54.02 34.79 42.35 37.51 49.74 42.35 27.41 37.33 31.80 39.78 32.25
40682 58.32 86.29 57.89 59.22 57.39 86.05 17.95 40.58 86.94 17.86 53.15
40705 43.49 37.92 0.73 30.63 47.90 46.28 31.38 10.71 -4.40 32.90 45.45
42261 62.01 90.39 60.13 58.15 59.78 56.78 55.10 49.09 77.20 52.53 85.15
42585 60.84 30.52 51.26 60.71 91.79 43.40 72.10 38.30 23.04 3.50 95.05
Mean 55.54 48.49 51.43 53.56 55.93 54.57 52.28 40.37 31.32 26.95 57.43
Mean-Rank 4.69 5.65 5.72 5.24 4.69 5.01 5.62 8.18 8.22 8.85 4.13
Top-3 11 12 6 5 15 11 9 1 6 4 21
Top-1 4 7 3 0 2 1 1 0 2 0 12
Table 10: Evaluation of clustering quality with the ARI metric on Synthetic Gaussian datasets (higher scores reflect better performance).
ID KM GMM AE-KM AE-GMM DEC IDEC IDC G-CEALS TabPFN SCARF ZEUS
0 86.16 82.13 83.44 79.66 87.06 77.26 60.64 58.57 65.42 4.06 87.65
1 88.66 87.36 81.13 79.61 85.84 75.62 44.44 54.10 40.26 4.21 72.04
2 81.30 26.46 71.57 78.04 86.65 78.91 53.53 48.59 20.42 11.44 82.96
3 95.02 36.77 92.03 76.44 92.74 89.81 65.03 62.67 50.57 22.29 95.76
4 90.64 92.97 82.65 77.80 91.61 84.51 77.56 69.70 61.65 8.80 92.85
5 89.83 25.85 86.52 69.06 90.01 88.31 70.38 67.70 20.51 10.31 90.72
6 96.21 98.07 85.99 94.26 94.24 92.39 80.18 73.15 86.50 6.51 97.89
7 92.59 30.52 81.73 79.06 91.35 89.80 46.44 58.10 26.62 10.14 93.60
8 88.29 89.62 81.15 76.53 87.56 78.37 72.72 50.30 68.54 13.06 84.00
9 93.36 98.54 68.58 85.42 91.14 85.29 76.64 56.89 65.02 11.37 94.55
10 79.26 93.32 72.75 90.90 72.54 73.51 77.64 48.47 89.15 -0.04 87.32
11 93.60 98.48 90.77 89.11 95.03 92.19 74.56 64.24 68.28 2.11 97.45
12 86.82 77.45 73.64 71.40 86.04 77.93 60.55 58.45 55.27 19.11 86.04
13 96.77 98.07 94.05 90.15 96.11 85.66 86.14 78.20 80.24 5.29 96.17
14 90.73 43.20 79.65 69.09 92.68 77.93 72.80 67.84 12.93 8.16 94.60
15 93.58 96.23 85.74 89.30 94.43 89.41 72.54 72.88 91.02 3.24 96.68
16 94.70 92.45 90.98 89.41 94.53 88.14 77.56 85.86 59.58 4.85 87.19
17 85.26 87.10 62.23 79.52 84.08 75.00 50.45 42.55 50.06 7.69 86.29
18 88.77 94.83 84.47 84.74 87.39 83.46 61.36 67.96 74.90 7.69 83.01
19 86.54 89.11 76.18 78.46 86.06 67.99 47.46 70.50 32.39 6.02 73.85
Mean 89.90 76.93 81.26 81.40 89.35 82.57 66.43 62.84 55.97 8.32 89.03
Mean-Rank 2.65 3.65 5.65 5.60 3.23 5.70 7.95 8.90 8.75 11.00 2.92
Top-3 16 13 0 1 15 0 0 0 1 0 14
Top-1 3 10 0 0 1 0 0 0 0 0 6
Table 11: Assessment of clustering quality based on the ARI metric on Synthetic Transformed datasets (higher score indicate better performance).
ID KM GMM AE-KM AE-GMM DEC IDEC IDC G-CEALS TabPFN SCARF ZEUS
0 60.56 79.33 46.31 68.98 76.93 46.35 56.67 29.42 6.36 0.94 87.42
1 93.30 96.55 76.57 83.99 91.29 70.91 66.59 73.09 16.59 1.22 89.40
2 89.76 96.31 57.79 65.99 94.42 66.34 72.25 48.34 10.00 1.10 95.88
3 70.16 50.38 56.23 76.06 83.06 60.09 65.13 54.54 15.36 1.99 80.98
4 44.32 80.63 47.60 44.77 68.39 50.12 36.00 49.37 19.14 1.48 73.66
5 91.49 96.51 74.48 89.11 86.46 77.84 60.00 66.47 15.12 1.25 81.63
6 60.98 81.63 61.97 76.29 75.58 61.89 69.43 69.84 18.83 2.11 87.02
7 76.80 83.98 71.37 77.69 86.56 77.49 76.75 56.51 18.94 3.20 89.57
8 89.77 77.10 85.53 84.42 93.33 92.96 90.69 78.87 24.26 5.91 98.11
9 84.40 98.48 82.64 93.03 97.14 81.02 91.02 65.62 9.01 1.66 97.20
10 82.47 91.91 61.00 86.16 87.58 46.08 70.42 39.27 14.11 0.46 86.42
11 90.77 99.32 81.41 91.18 94.78 88.12 80.75 70.29 18.31 1.72 93.19
12 67.64 91.80 52.87 63.56 85.45 60.61 73.66 55.59 28.50 3.21 87.75
13 85.99 50.80 45.48 35.34 29.04 12.91 30.22 32.89 35.78 6.01 91.60
14 93.52 54.56 15.81 22.79 94.12 38.21 68.54 15.10 15.08 5.54 98.49
15 36.62 30.93 37.71 43.70 42.78 40.65 32.64 29.70 16.68 0.63 28.96
16 95.62 65.11 90.13 94.97 97.47 95.32 84.79 98.27 17.27 9.58 99.01
17 60.54 95.74 56.24 91.71 78.72 59.11 82.54 13.29 7.33 -0.10 96.44
18 63.92 17.39 58.47 70.21 54.61 47.93 64.15 10.01 3.79 0.65 79.70
19 62.12 79.06 49.36 65.77 81.07 51.19 63.41 26.84 2.73 1.06 84.13
Mean 75.04 75.88 60.45 71.29 79.94 61.26 66.78 49.17 15.66 2.48 86.33
Mean-Rank 4.80 3.50 6.85 4.50 3.20 6.35 6.05 7.80 9.75 11.00 2.20
Top-3 4 14 0 6 15 2 1 1 0 0 17
Top-1 0 8 0 1 1 0 0 0 0 0 10
@@ F.2 Are ZEUS’s assignments well calibrated?
Tables 12, 13 and 14 contain detailed extension of Table 3 from the main part of the paper. Similar to the Tables in Appendix F.1, these also include additional rows: Mean, Mean-Rank, Top-3, and Top-1, which aggregate the results presented in each table.
ZEUS consistently attains at least the second position across all statistics in every table presented in this section. For the Synthetic Transformed datasets, it is undeniably the best. However, for the OpenML datasets, it is outperformed by IDEC, and for the Synthetic Gaussian datasets, GMM takes the lead in the Top-1 metric, while for the other statistics, k-means performs better.
Table 12: Soft clustering performance of ZEUS versus competing methods on real-world datasets, measured by Brier score (lower is better).
@@
Table 13: Soft clustering performance (Brier score) of ZEUS compared to baseline methods on Synthetic Gaussian datasets (lower score reflect better quality).
@@
Table 14: Evaluation of soft clustering quality on Synthetic Transformed datasets using Brier score (lower score indicate better performance).